OpenStudy (anonymous):

If g(x) has a vertical translation down of 3 units of f(x)=2x-1 whats is the rule od g(x)

OpenStudy (anonymous):

So we need to subtract 3 from the -1 on f(x) f(x) = 2x - 1 g(x) = 2x - 1 - 3 g(x) = 2x - 4 or g(x) = 2(x - 2)